OpenStudy (anonymous):

Rewrite the following as an algebraic expression of x. csc(arccos(x)) State the domain on which the equivalence is valid.

OpenStudy (anonymous):

Using the pythagorean identity: \[\csc(\cos^{-1} (x))=\frac{ 1 }{ \sin(\cos^{-1} (x)) }=\pm\frac{ 1 }{ \sqrt{1-\cos ^{2}(\cos^{-1}(x))} }=\pm \frac{ 1 }{ \sqrt{1-x^{2}} }\] Domain: \[\forall x \in \mathbb{R} \]

OpenStudy (anonymous):

thanks